OpenStudy (anonymous):

How much pure acid must be added to 50 mL of a 25% acid solution to product a mixture that is 60% acid?

OpenStudy (anonymous):

at the moment you have \(25\%\) of \(50\) or \(12.5\) mL of acid when you add \(x\) liters of acid, you will have \(12.5+x\) liters of acid total, and \(50+x\) liters of solution which you want to be \(60\%\) acis since you want them to be equal, set \[12.5+x=.6(50+x)\] and solve for \(x\) which might be easier if you multiply by ten and solve \[125+10x=6(50+x)\]